OpenStudy (anonymous):

Help!?! 3. A concert orchestra plays a concert at your school, producing a sound intensity of 90 dB. At a point in the performance, a soloist takes over and plays with a sound intensity of 70 dB. A. How many times more intense is the entire orchestra than the soloist? Show your work. B.During the performance, someone's cell phone rings in the audience at a volume of 60 dB. How many times louder is the orchestra than the sound of the cell phone ringing? Show your work. C. A few weeks later, you attend a rock concert at which the band plays at 110 dB. How many time

OpenStudy (anonymous):

Loudness is expressed in a logarithmic scale. The unit that you get from the log of the intensity is bel. So, if a given sound is 100 times louder, it gives log (10^2) = 2. So, it is 2 bel. However, the decibell is the accepted unit, so you multiply the value by 10 and you get 20 dB.

OpenStudy (anonymous):

On your task, A. The difference is 90 - 70 = 20 dB. 20 dB = 2 bel, so 10^2 = 100 times louder. B. Similarly, 90 - 60 = 30 dB, so 10^3 = 1000 times louder. C. A part of the task is missing but I'm sure you get the idea, so that you can do it yourself.

OpenStudy (anonymous):

Thank you.